Required information Yvonnes father was a true believer in giving back. He endowed a program 35 years ago to help students
Answer #1

Answer:$5184.73

If the sum is able to provide perpetual income of $22,000 the present value of the sum will be:

PV of perpetuity=PMT/i=22000/11%=$200,000

Since the sum is 200,000 now its present value 35 years back shall be:

Present Value (P)=C/{[(1+r)^n]}
Where,
C=Future Value= 200000
r=ROI= 0.11
n=No. of periods= 35
Now putting values in formula,
= 200000/{[(1+ 0.11)^35 ]}
= 200000/{[( 1.11)^35 ]}
= 200000/{[ 38.57485 ]}
= 5184.7251

If helpful, Thumbs UP please:)